solve using the substitution method
1/3x+1/5y=7, 1/6x-2/5y=-4

Answer by Alan3354(69443)

1/3x+1/5y=7
1/6x-2/5y=-4 times 2 --> x/3 - 4y/5 = -8 --> x/3 = 4y/5 - 8
----------
x/3 + y/5 = 7
sub for x/3
(4y/5 - 8) + y/5 = 7
4y - 40 + y = 35
5y = 75
y = 15
-----------
x/3 + y/5 = 7
x/3 + 3 = 7
x/3 = 4
x = 12